OpenStudy (anonymous):

factor completely: 6x^2-5x+1

OpenStudy (anonymous):

This one is tougher... I'll tell you my thought process

OpenStudy (anonymous):

the last term's factors usually add up to the center term. Since the only factors are 1 and 1 and they don't add up to 5, I look then at the first term

OpenStudy (anonymous):

i have to show my work luis

OpenStudy (anonymous):

you only have two choices if it is going to factor using integers either \((2x-1)(3x-1)\) or \((6x-1)(x-1)\) you can check which one works

OpenStudy (anonymous):

i know the answer is (2x-1)(3x-1) but how do we get to that?

OpenStudy (anonymous):

the only way to get the 1 at then end \(1\times 1\) or \(-1\times -1\) and since the middle term is negative it must be \(-1\times -1\) that is why there are only two options, as now the only way to get \(6x^2\) is \(3x\times 2x\) or \(6x\times x\)